OpenStudy (sugarmochi):

please help solve 32x^2-40x-12=-8x+12

OpenStudy (will.h):

set the entire expression equal to zero try and solve that out for now and tag me when you done that

OpenStudy (sugarmochi):

\[32x^2-38x-24=0 \]

OpenStudy (sugarmochi):

do i start from this? @Will.H

OpenStudy (mathmale):

Starting from "32x^2-40x-12=-8x+12," manipulate the equation so that all terms are on the left side. The right side will automatically be 0 if you'd do that. Simplify the left side as much as possible. You should get 3 terms. Write them in descending order by powers of x. Your result?

OpenStudy (sugarmochi):

32x^2-38x-24=0 i think..

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

i think so too

OpenStudy (misty1212):

oh actually there is an arithmetic mistake there

OpenStudy (misty1212):

\[-40+8=-32\]

OpenStudy (misty1212):

start with \[32x^2-32x-24=0 \]

OpenStudy (sugarmochi):

oh sorry bout the mistake !

OpenStudy (misty1212):

then divide by the greatest common factor first

OpenStudy (misty1212):

if that is not clear, divide by 8 , since 8 goes in to each term evenly and get\[4x^2-4x+3=0\] which, by some miracle, factors

OpenStudy (will.h):

or use the famous quadratic formula \[\frac{ -b \pm \sqrt{b^2 -4ac} }{ 2a }\]